3. 0/4 points Previous Answers An FM station broadcasts classical music at 88.8 MHz (megahertz, or 106 Hz). Find the waveleng need answer to 3 & 4 please
Answer #1

3.

We have

Frequency = 88.8 MHz = 88.8 \times 106 Hz

Now the formula of wavelength

wavelength = Speed of light / frequency

where speed of light c = 3.0 \times 108 m/s

so

Wavelength \lambda = 3.0 \times 108 m/s / 88.8 \times 106 Hz = 3.78 m

to convert wavelength from m to nm

we know that

1 m = 1 \times 109 nm

so

Wavelength = 3.78 \times 109 nm

Also

1 m = 1 \times 1010 A0   

then

Wavelength = 3.78 \times 1010 A0

4.

we have

wavelength = \lambda = 1.6 A0 = 1.6 \times 10-10 m

Energy E = hc / \lambda

where h = Plancks constant = 6.626 \times 10-34 J.s

c = speed of light = 3.0 \times 108 m/s

So

Energy = 6.626 \times 10-34 J.s \times 3.0 \times 108 m/s / 1.6 \times 10-10 m

= 12.42 \times 10-16 J